Pebbles are poured out of a tube at one cubic meter per second. It forms a pile which has the shape of a cone. The height of the cone is equal to the radius of circular base. How fast is the pile of pebbles rising when it is 2 meters high


1
Expert's answer
2021-12-01T17:51:59-0500

Let V = volume of pebble in pile at a time t


"Given\ \frac{dv}{dt}=1m^3/s"


"Find \ \frac{dh}{dt}, when \ h=2m"


"V=\frac{1}{3}\pi r^2h"


Diameter = height

2r = h

"r=\frac{h}{2}"


"V=\frac{1}{3}\pi (\frac{h}{2})^2h"


"V=\frac{1}{12}\pi h^3"


"\frac{dv}{dt}=\frac{1}{12\pi}(3h^2.\frac{dh}{dt})"


"\frac{dv}{dt}=\frac{3}{12}\pi h^2 \frac{dh}{dt}"


because "\frac{dv}{dt}=1"


"\frac{3}{12}\pi h^2 \frac{dh}{dt}=1"


"\frac{dh}{dt}=\frac{12}{3\pi h^2 }"


"\frac{dh}{dt}|_{h=2}=\frac{12}{3\pi (2)^2 }"


"=0.3183m\/s"

The pile pebbles rise at "0.3183m\/s"